OpenStudy (anonymous):

What is the slope of the line passing through the points (2, 5) and (0, −4)?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

\[\frac{-4-5}{0-2}=\frac{9}{2} \]

OpenStudy (skullpatrol):

$$\Huge \text{slope = } \frac{rise}{run}$$

OpenStudy (anonymous):

this stuff hella confuses me uhhhh

OpenStudy (anonymous):

tru